Systems Staff Meeting Notes 5/17/99

Systems Staff Meeting Notes
May 17, 1999

Present: Rich Bennett, Brian Burton, Christian Conrad, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng, Suzy Shaw, Ying Tang
  1. Debra received a strange laptop checkout email from Germany. They were very precise about what they needed - and the request was urgent!
  2. Staff Locator (PAULA) - please be as precise as possible with the details of your whereabouts. Need more information than just a building name.
  3. Keys & locks - we're in the process of getting Intellikey in Systems. All of Systems doors as well as room 219 closet & the LAC equipment drop off room will be converted.
Review Minutes of 5/10/99
  1. RIR #1 - the 40 Cows have all been distributed.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Even more HD's are on order - we have 4 incoming & at least 5 requests. We're only upgrading the 133's & up - the rest will be replaced in the next fiscal year.
  2. One more large Gateway order will be made this fiscal year.
  1. Basic training program to start in August. It will be a three-day event, to be conducted by an Orlando firm and should be attended by all of the liaisons.
  1. Web LUIS outage Sunday morning - the Library page was accessible, but not the catalogs. NERDC was doing a mainframe upgrade.
  2. Web LUIS outage led to a COWS1 password lock out. Bill has worked on the permissions so that after 10 minutes the password attempts counter resets. If there are more than 7 attempts to log-on, it will lock out. Within 10 minutes (if left alone) it will unlock. We may have to have generic sign-ons for each building if this 10 minute reset does not help the problem.
  3. The campus-wide video broadcast, "The Shape of the River", went reasonably well last week. This was the first large-scale attempt and we'll probably attempt to do more in the future as the quality improves. The maximum data stream for this broadcast was 80K.
  4. Wiring - for the next couple of weeks, in the early morning & late afternoon hours, Ashley & crew will be working to clean up the wiring so it's up to standards in 219 West & MSL.
  5. The cabinets in Music, AFA & Education have been installed. This should lessen the noise in those areas. After finishing the encumbrance on those this week, we will be able to place the new Gateway order. Barbara Oliver has informed Ashley that the closing date for blanket accounts is May 27.
  6. Vault inventory/shelf updates - removal of equipment or hardware from the vaults should be reported on an inventory list that is hanging from the shelves in the vault. Everett & Bobbie will update the spreadsheet with the information. This does not include the supply shelf.
Software Projects Status
  1. Student Patron Update for Summer A is complete.
  2. IR database is looking good. A Microsoft Access module may be added for SQL server pointers to OBDC. It can be read/write or read-only. This will allow for specific queries without the hard code.
DP Coordinator
  1. The latest batch of yearly Compustat tapes will not copy. They are producing the same error messages as all of the other tapes. Bill talked to Andy Halula and has been given the name of an individual who can read the IBM binary. We will try swapping a couple of tapes with them. If we can read their tapes, then it would seem to indicate that the tapes that Compustat has sent to us are not formatted correctly. If we are unable to read their tapes, then it would seem to indicate a problem on NERDC's end.
  1. Everett out today.
  2. Bill & Suzy out May 24-28.
  3. Brian out May 28-June 1.
  4. Brian out June 14-18 at an InterDev Conference.
  5. Winston out July 1 & 2.
  6. Bobbie and Winston out July 12-16.

Back to the Systems Homepage

Last updated February 18, 2000
by Debra Harris